# Optimal Planning of Electric Vehicle Charging Stations: Integrating Public Charging Networks and Transportation Congestion

### v1

@@ Abstract
The transition to electric vehicles (EVs) marks a pivotal shift in personal mobility, driven by policy incentives and automotive innovations. However, the expansion of EVs for long-distance travel is hindered by charging time concerns, the sparse distribution of charging stations, and the worsening waiting times due to congestion. The main objective of this work is two-fold: 1) first, to comprehensively analyze the existing public charging station robustness and effectively strategize for the new ones, and 2) secondly, to select the optimal chargers for long-distance journeys, by estimating the waiting time from current traffic congestion. This is achieved by accompanying effective EV charging strategies, pinpointing on the congestion points from the existing traffic, and the robustness of the current charging station infrastructure. Utilizing a real-time transportation and charging station dataset in Texas, we identify optimal charger placement strategies to minimize travel time by examining the congestion and charging time trade-offs. Our findings suggest that maximizing the constant current phase during charging enhances efficiency, crucial for long-distance travel. On the contrary, we also explore the negative impact of congestion on travel times and we conclude that sometimes it might be beneficial to exceed the constant current phase to avoid the congested charging stations.
Index Terms:
EV Charging Network Planning, Network Optimization, Traffic Congestion, Charging Strategy
@@ I Introduction
The transition to electric vehicles (EVs) marks a pivotal shift in personal transportation, motivated by the urgent need to mitigate carbon emissions and reduce reliance on fossil fuels. As EVs emerge as the standard for daily transportation, the critical challenge lies in enabling efficient long-distance EV travel, highlighting the importance of a comprehensive and accessible charging infrastructure. The proliferation of public charging stations is a positive trend, yet optimizing their distribution is key to enhancing the robustness and reach of the charging network for extended journeys.
@@
This paper examines the interplay between the availability of public chargers with their respective waiting times according to current traffic, the characteristics of EV batteries, and strategies for minimizing their charging periods. Determining the optimal placement and capacity of charging stations not only enhances user experience but also mitigates adverse effects on the distribution network at a reduced cost [1]. Archana and Rajeev [2] introduce a reliability index for allocating EV charging stations within a 33-bus distribution network, while [3, 4, 5] integrate EV charging station allocation with renewable generation in distribution networks. Locations of EV charging stations are decided in [6, 7, 8], where traffic flows are taken care of. However, the existing studies have not accounted for real traffic networks. In this research, we focus specifically on planning for the Dallas-Fort Worth (DFW) area using a real road-network and existing traffic.
@@
The Constant Current Constant Voltage (CCCV) technique is frequently employed for charging. Introduced by Shi et al. [9], a pioneering remaining charging time estimation algorithm emphasizes the significance of the constant current (CC) phase to uphold high charging efficiency for long-haul travel. As EVs transition to the constant voltage (CV) phase, where voltage is steady but current diminishes, the overall charging power drops, prolonging charging duration and decreasing efficiency. Considering the diverse battery characteristics of EVs, the CC phase correlates with specific driving ranges, identifying the optimal charging point as the juncture where cell voltages hit their peak operational level. Distinct EV models feature varied battery capacities and preferred charging State of Charge (SOC) ranges, affecting the determination of optimal charging site locations. For instance, the ideal recharging stops for a Tesla differ from those for a GM Bolt, even if the destination remains the same. Route planning involves finding route based on criteria such as driving distance, travel time, or energy consumption. Dijkstra’s algorithm is widely used for this purpose [10], although it can be slow for large graphs. Techniques like the A* algorithm use heuristics for faster searches [11], while contraction hierarchies [12] add shortcuts to speed up path finding in preprocessing steps.
@@
Route planning for electric vehicles introduces challenges due to battery constraints, including limited range and recuperation. This leads to a constrained shortest path problem [13], where finding the shortest path considering battery constraints is essential. Modified versions of the Dijkstra’s algorithm or A* can calculate all Pareto optimal paths, although this is more complex. Contractions hierarchies can also be used to speed-up multi-criteria path finding and to solve the shortest path in acceptable time [14].
@@
Baum et al. [15] proposed a method for solving the electric vehicle constrained shortest path problem, prioritizing paths that comply with battery constraints while minimizing travel time. Their approach suggests driving speeds and avoids the need to compute all Pareto optimal paths. They accelerate queries using a combination of contraction hierarchies and the A* algorithm.
Razo and Jacobsen [16] proposed a model that relies on mobile network reservations to estimate charging station waiting times, aiding vehicle selection. Gusrialdi et al. [17] introduced a model using vehicle-to-vehicle communication to decide charging station stops. Tan and Wang [18] utilized game theory to optimize travel costs for users and maximize revenues for charging stations, considering freely set hourly prices and grid overloading. Yang et al. [19] developed a cooperative game theory approach to enhance convergence speed and robustness.
@@
To enhance practicality, our work in this veins estimates the shortest path by considering the actual traffic and real road network of the DFW area. By utilizing real-time traffic data and congestion analysis, we can accurately forecast waiting times. With the increase in EV adoption, a rise in vehicles on the road leads to inevitable congestion at charging sites, especially near major thoroughfares. This paper introduces an optimal charging strategy that accommodates congestion considerations and the efficiency trade-offs of extended CV phase duration to enhance total travel time. By aligning charging stations with driving ranges tied to optimal charging efficiency phases, the study aims to boost the practicality and appeal of EVs for long-distance travel.
@@
The rest of the paper is organized as follows. Section II discusses our methodology in detail. It begins with our strategy for future charger locations. Next, we discuss an algorithm to select appropriate charging stations for minimizing the total travel time and charging times. Section III delves into a case study assessing the performance of our algorithm using real dataset, integrating the densely populated Dallas-Fort Worth area with traffic congestion points, charging stations, and the actual road network. The paper concludes in Section IV, summarizing the findings and their implications for future research and policy-making.
@@ II Model For EV Charger Network Planning and EV Long-distance Planning
Focusing on charger locations in Texas, sourced from the Department of Energy (DOE) database [20], our objective is to devise a customized charger network that identifies optimal charger locations for various EV models. We utilize vehicle transportation dataset from the Texas transportation website [21], specifically annual average 24-hour vehicle counts across Texas, with a current EV percentage of 1.4% of the entire traffic [22]. The selection of optimal charger locations relies on a comprehensive understanding of travel distances, duration, congestion levels, and the Constant Current Constant Voltage (CCCV) charging method, specifically considering the lithium-ion battery’s most efficient SOC window during the CC phase [23]. Given the diverse battery characteristics across EVs, the CC phase is associated with specific driving ranges, with the optimal charging point occurring when cell voltages reach their maximum.
@@ II-A Current Charger Network Analysis
Percolation analysis demonstrated the network’s robustness over graphs [25] to random failures but revealed the vulnerabilities to targeted disruptions of key nodes. This insight suggests enhancing critical nodes with redundancy and rapid-response repair capabilities to ensure continued network functionality.
@@ II-C Selecting Prime Charging Spots for Long-Haul EV Travel
$\sum_{i\in V}x_{ij}=\sum_{i\in V}x_{ji}$ for $\text{for }j\in N,\{(i,j),(j,i)\}\in E$ (2)
if $\text{if }x_{ij}=1\ \Rightarrow\ y_{i}=1\text{ and }y_{j}=1$
[…]
@@ III Case Study for EV Charger Network
Using the Python libraries NetworkX [28], OSMnx[29], and Pandas[30], we developed a detailed and realistic model of the DFW road network. This model incorporates comprehensive traffic data and geographic information, enabling dynamic adjustments to route planning based on current traffic conditions. The core of our approach is the development of a weighted graph, where nodes represent intersections or points of interest (such as EV charging stations), and edges symbolize the roadways connecting these points. Weights assigned to each edge correspond to travel times, which are adjusted dynamically based on data-driven predictions of traffic congestion and anticipated wait times at EV charging stations. This predictive capability is crucial for real-time traffic management, allowing the system to adapt routes on the fly to minimize delays and improve overall travel efficiency. The travel planning between two distant points, obtained by Algorithm 1 is shown in Fig. 5.

### v2

@@ Abstract
The adoption of electric vehicles (EVs) represents a critical shift in personal mobility, fueled by policy support and advancements in automotive technology. However, the expansion of EVs for long-distance travel is hindered by charging time concerns, the sparse distribution of charging stations, and the worsening waiting times due to congestion. The main objective of this work is two-fold: 1) first, to comprehensively analyze the existing public charging station robustness and effectively strategize for the new ones, and 2) secondly, to select the optimal chargers for long-distance journeys, by estimating the waiting time from current traffic congestion. This is achieved by accompanying effective EV charging strategies, pinpointing on the congestion points from the existing traffic, and the robustness of the current charging station infrastructure. Utilizing a real-time transportation and charging station dataset in Texas, we identify optimal charger placement strategies to minimize travel time by examining the congestion and charging time trade-offs. Our findings suggest that maximizing the constant current phase during charging enhances efficiency, crucial for long-distance travel. On the contrary, we also explore the negative impact of congestion on travel times and we conclude that sometimes it might be beneficial to exceed the constant current phase to avoid the congested charging stations.
@@ I Introduction
A continuous shift to electric vehicles (EVs) represents a pivotal transformation in personal transportation, driven by the pressing need to lower carbon emissions and decrease dependence on fossil fuels. As EVs become the norm for daily transportation, the critical challenge lies in enabling efficient long-distance EV travel, highlighting the importance of a comprehensive and accessible charging infrastructure. The proliferation of public charging stations is a positive trend, yet optimizing their distribution is key to enhancing the robustness and reach of the charging network for extended journeys.
@@
This paper examines the interplay between the availability of public chargers with their respective waiting times according to current traffic, considering the characteristics of EV batteries, and strategies for minimizing their charging periods. Determining the optimal placement and capacity of charging stations is important for enhancing user experience while minimizing overall charger investment costs. Authors in [1] introduce a reliability index for allocating EV charging stations within a 33-bus distribution network. Location planning of EV charging stations have been studied in [2, 3, 4], considering a generic traffic, however, the existing studies have not accounted for real traffic networks. In this research, we focus specifically on considering a real transportation network along with the existing real traffic flow there.
@@
A Constant Current Constant Voltage (CCCV) technique is frequently employed for charging. Introduced by Shi et al. [5], a pioneering remaining charging time estimation algorithm emphasizes the significance of the constant current (CC) phase to uphold high charging efficiency for long-haul travel. Distinct EV models feature varied battery capacities and preferred charging State of Charge (SOC) ranges, affecting the determination of optimal charging site locations. For instance, the ideal recharging stops for a Tesla differ from those for a GM Bolt, even if the destination remains the same.
Route planning involves finding route based on criteria such as driving distance, travel time, or energy consumption. Dijkstra’s algorithm is widely used for finding the shortest distance in a network[6], although it can be slow for large graphs. Techniques such as A* algorithm use heuristics for faster searches [7].
@@
Route planning for electric vehicles becomes more complex due to battery constraints, including limited range and energy recuperation capabilities. This introduces the constrained shortest path problem [8], where finding the shortest path considering battery constraints is essential. Modified versions of the Dijkstra’s algorithm or A* can calculate all Pareto optimal paths, although this increases computational complexity. Contractions hierarchies can also be used to speed-up multi-criteria path finding and to solve the shortest path in acceptable time [9].
@@
Baum et al. [10] proposed a method for solving the electric vehicle constrained shortest path problem, prioritizing paths that comply with battery constraints while minimizing travel time. Their approach suggests driving speeds and avoids the need to compute all Pareto optimal paths. They accelerate queries using a combination of contraction hierarchies and the A* algorithm. Razo and Jacobsen [11] proposed a model that relies on mobile network reservations to estimate charging station waiting times, aiding vehicle selection. Tan and Wang [12] utilized game theory to optimize travel costs for users and maximize revenues for charging stations, considering freely set hourly prices and grid overloading.
@@
To enhance practicality, our work in this veins estimates the shortest path by considering the actual traffic and real road network of the Dallas Fort Worth (DFW) area. By utilizing real-time traffic data and congestion analysis, we can accurately forecast waiting times. With the increase in EV adoption, a rise in vehicles on the road leads to inevitable congestion at charging sites, especially near major thoroughfares. This paper introduces an optimal charging strategy that accommodates congestion considerations and the efficiency trade-offs of extended CV phase duration to enhance total travel time. By aligning charging stations with driving ranges tied to optimal charging efficiency phases, the study aims to boost the practicality and appeal of EVs for long-distance travel.
@@
The rest of the paper is organized as follows. Section II discusses our methodology in detail. It begins with a comprehensive analysis of current traffic conditions and the distribution of EV charging stations across Texas, identifying key areas with insufficient charging infrastructure, and provides recommendations for future charger locations. Next, we introduce an algorithm to select appropriate charging stations for minimizing the total travel time and charging duration. Section III delves into a case study assessing the performance of our algorithm using real dataset, integrating the densely populated Dallas-Fort Worth area with traffic congestion points, charging stations, and the actual road network. The paper concludes in Section IV, summarizing the findings and their implications for future research and policy-making.
@@ II Model For EV Charger Network Planning and EV Long-distance Planning
Focusing on charger locations in Texas, sourced from the Department of Energy (DOE) database [13], our objective is to devise a customized charger network that identifies optimal charger locations for various EV models. We utilize vehicle transportation dataset from the Texas transportation website [14], specifically annual average 24-hour vehicle counts across Texas, with a current EV percentage of 1.4% of the entire traffic [15]. The selection of optimal charger locations relies on a comprehensive understanding of travel distances, duration, congestion levels, and the Constant Current Constant Voltage (CCCV) charging method, specifically considering the lithium-ion battery’s most efficient SOC window during the CC phase [16]. As EVs transition to the constant voltage (CV) phase, where voltage is steady but current diminishes, the overall charging power drops, prolonging charging duration and decreasing efficiency. Considering the diverse battery characteristics of EVs, the CC phase is associated with specific driving ranges, with the optimal charging point occurring when cell voltages reach their maximum.
@@ II-A Current Charger Network Analysis
Percolation analysis demonstrated the network’s robustness over graphs [18] to random failures, but it also exposed vulnerabilities in the event of targeted attacks on critical nodes. This insight suggests enhancing critical nodes with redundancy and rapid-response repair capabilities to ensure continued network functionality.
@@ II-C Selecting Prime Charging Spots for Long-Haul EV Travel
$\sum_{i\in V}x_{ij}=\sum_{i\in V}x_{ji}$ $\hskip-102.43008pt\text{for }j\in N,\{(i,j),(j,i)\}\in E$ (2)
$\text{if }x_{ij}=1\ \Rightarrow\ y_{i}=1\text{ and }y_{j}=1$
 $\hskip 51.21504pt\text{ for }i/j\in V_{c}\text{ and }(i,j)\in E$ (3)
$c_{ij}x_{ij}\leq\lambda_{\text{threshold}}$  $\hskip-51.21504pt\text{ for }(i,j)\in E$ (4)
$x_{ij}\in\{0,1\}$  $\hskip-51.21504pt\text{ for }(i,j)\in E$ (5)
$y_{i}\in\{0,1\}$  $\hskip-51.21504pt\text{ for }i\in V_{c}$ (6)
@@ III Case Study for EV Charger Network
Using the Python libraries NetworkX [21] and OSMnx[22], we developed a detailed and realistic model of the DFW road network. This model incorporates comprehensive traffic data and geographic information, enabling dynamic adjustments to route planning based on current traffic conditions. The core of our approach is the development of a weighted graph, where nodes represent intersections or points of interest (such as EV charging stations), and edges symbolize the roadways connecting these points. Weights assigned to each edge correspond to travel times, which are adjusted dynamically based on data-driven predictions of traffic congestion and anticipated wait times at EV charging stations. This predictive capability is crucial for real-time traffic management, allowing the system to adapt routes on the fly to minimize delays and improve overall travel efficiency. The travel planning between two distant points, obtained by Algorithm 1 is shown in Fig. 5.